OpenStudy (anonymous):

solve the differential equation- [4e^(-y)*sinx-1)dx-dy=0

OpenStudy (anonymous):

[4(e^-y)*sinx-1)dx=dy

OpenStudy (anonymous):

dy/dx=[4(e^-y)*sinx-1]

OpenStudy (anonymous):

after dat how should i seperate?

OpenStudy (experimentx):

d(x+y)/dx = e^(-y)sin(x) let x+y=z, y=z-x

OpenStudy (anonymous):

frm where does(x+y) comes?

OpenStudy (anonymous):

ok u took 1 to the other side

OpenStudy (experimentx):

yes

OpenStudy (experimentx):

use separation of variables after that

OpenStudy (anonymous):

ok m trying it out

OpenStudy (anonymous):

and cant we solve it by bernoullis eqn?

OpenStudy (experimentx):

IDK ... haven't tried. my guess is no

OpenStudy (anonymous):

m not able to solve it by ur method! :/

OpenStudy (experimentx):

d(x+y)/dx = e^(-y)sin(x) => dz/dx = e^(x-z) sin(x) => e^z dz = e^x sin(x) dx finish him off

OpenStudy (anonymous):

could u plsz show me the steps so dat i could recognize my mistake

OpenStudy (experimentx):

and 4 is missing!!

OpenStudy (anonymous):

ohhhh!!!! ok m solving it

OpenStudy (anonymous):

can u plz tell me the integration of e^x8cos x?

OpenStudy (anonymous):

should i take cos x as u and e^x as v?

OpenStudy (experimentx):

there are 3 ways of integrating it Integrating by parts changing sin(x) = e^(ix) and taking Imaginary values or use symmetricity http://www.wolframalpha.com/input/?i=Integrate%5BSin%5Bx%5D+E%5Ex%2C+x%5D

OpenStudy (anonymous):

pls tell me u*v metho0d

OpenStudy (experimentx):

\[ \frac{d}{dx} (e^{x} \sin(x)) = e^x \sin (x)+e^x \cos (x) \\ \frac{d}{dx} (e^{x} \cos(x)) = e^x \cos (x)-e^x \sin (x) \] subtract second from first .. and integrate on both sides

OpenStudy (anonymous):

its not gettin cancelled

OpenStudy (experimentx):

the left side is \[ \frac{d}{dx} (e^x\sin(x) - e^x \cos(x)) = 2 e^x \sin(x)\] integrate both sides and get your result.

OpenStudy (experimentx):

\[ \frac 1 2\int \frac{d}{dx} (e^x\sin(x) - e^x \cos(x)) dx = \int e^x \sin(x)dx\]

OpenStudy (anonymous):

ok thanku for ur help!

OpenStudy (experimentx):

you are welcome